OpenStudy (anonymous):

The area of a triangle is 16 square units. The base of the triangle is x+4. The height of the triangle is x. Find x.

OpenStudy (anonymous):

16=x+4*x*.5 16=(x^2+4x)/2 32=x^2+4x I don't know what to do next...

OpenStudy (anonymous):

okay thanks!(:

OpenStudy (anonymous):

quadratic formula

OpenStudy (anonymous):

there you go! the quadratic formula

OpenStudy (anonymous):

http://www.mathsisfun.com/quadratic-equation-solver.html it's -8 and 4

OpenStudy (anonymous):

aha thanks